OpenStudy (anonymous):

What is the value for x in the proportion x over x plus ten equals seven over twelve ?

sam (.sam.):

\[\frac{x}{x+10}=\frac{7}{12}\]

OpenStudy (anonymous):

so what stands for x ?

sam (.sam.):

14

OpenStudy (anonymous):

x/ (x+10) =7/ 12 Cross product: 12x = 7x + 70 ->5x = 70 => x = 70/5 = 14